OpenStudy (anonymous):

Solve the following if x = 3 & y = 4: 2x + y =

OpenStudy (anonymous):

Plug it in! 2(3) + 4 = 6 + 4 = 10 =)

OpenStudy (anonymous):

do i solve it to???

OpenStudy (anonymous):

Just substitute 3 in every time you see x and 4 in every time you see y in your expression.

OpenStudy (anonymous):

What do you mean you solve it? =/ You just have to plug the values in and find out what your equation equals too =)

OpenStudy (anonymous):

were did you get 6 from???

OpenStudy (anonymous):

2 times 3

OpenStudy (anonymous):

It says 2x that means 2 times x. If x = 3 then that is 2 times 3 = 6

OpenStudy (anonymous):

You mulitply 2(3) as the equation said to mulitply 2(x) as x = 3 So then it would be 2(3) = 6

OpenStudy (anonymous):

omg this is so confusing!!

OpenStudy (anonymous):

Look Wiz..Your not opening up your mind, your frustrating in the first place..Lets take this step by step Your are GIVEN that x= 3 and y = 4 Your also GIVEN the equation you need to insert these values..That equation is: 2x + y We are supposed to figure out what this expression will EQUAL when you plug in the values for x and y So we know that x = 3 and y = 4 YOU PLUG THEM IN: 2(3) + 4 Multiply 2 times 3 = 6 and add 4 to 6 = 10

OpenStudy (anonymous):

ohh thank you ur goood!!!!!

OpenStudy (anonymous):

Anytime Wiz =) Just dont stress